OpenStudy (anonymous):

Find the equation to the line tangent f(x)= sq(x+3) at (1,2).. Did i do it correctly??

OpenStudy (anonymous):

Slope m= y2-y1/x2-x1 m= 3-2/6-1 (1,2)(6,3) m=1/5 equation: y-2=1/5(x-1)